OpenStudy (anonymous):

suppose, f(x+h)-f(x)/h = -2h(x+4)^2-h^2/h(x+h+4)^2(x+4)^2 Find the slope of the tangent line at x=3

OpenStudy (anonymous):

\[\frac{ f(x+h)-f(x) }{ h } = \frac{ -2h(x+4)^2-h^2 }{ h(x+h+4)^2(x+4)^2 }\]

OpenStudy (anonymous):

is this suppose to be a limit? h->0

OpenStudy (anonymous):

It's finding the slope of a tangent line at x=3.

OpenStudy (anonymous):

yes i know the first part is normally written \[\lim_{h \rightarrow 0} \frac{f(x+h)-f(x)}{h}\] which is why im wondering whether or not you're missing the limit

OpenStudy (anonymous):

they do not include the on the question, i know what you mean though

OpenStudy (anonymous):

in that case, just divide out the h substitute x=3 and leave all the remaining h's and if you can simplify, then do so

OpenStudy (anonymous):

\[-\frac{ 2x+h+8 }{ (x+4)^2(x+h+4)^2 }\] is what i get when i simplify

OpenStudy (anonymous):

you're solving for the slope at x=3 so substitute 3 for x

OpenStudy (anonymous):

\[\frac{ 14h }{ (49)(49h) }\]this is what i get when i plug in 3 for x

OpenStudy (anonymous):

you made an error in calculations when doing the denominator (x+h+4)^2 (3+h+4)^2 (7+h)^2 49+14h+h^2 adding is not the same as multiplication